# Force-free wave interaction in magnetar magnetospheres: Computational modeling in axisymmetry

### v1

@@ 3.3 Prompt generation of FMS waves
Figure 5: As Figure LABEL:fig:a1bounce but for a single Alfvén wave injected with $b_{\rm eq}=0.39$ and half a wavelength ($cT/\lambda=0.5$, $\lambda=R_{\ast}/2$) at $\theta_{s}=45^{\circ}$ with width $\Delta\theta=20^{\circ}$ and resolution $N_{r}(N_{R_{\ast}})\times N_{\theta}=19968(512)\times 1600$. The fast wave amplitude relative to the magnetic field can develop extended electrically dominated zones with $E>B$ (green contours, panel d). We provide an animated version of this figure as supplementary material (Mahlmann et al., 2024b).
@@
Figure 6: As Figure 5, but for two Alfvén waves launched symmetrically and out-of-phase ($\varphi=\pi$) in both hemispheres. The Alfvén waves interact in the equatorial region (panels B/C/D) and alter the fast wave generation. We provide an animated version of this figure as supplementary material (Mahlmann et al., 2024c).
@@
Figure 7: Conversion efficiency of injected Alfvén wave energy to outgoing FMS wave energy using $N_{r}(N_{R_{\ast}})\times N_{\theta}=4992(128)\times 400$. Panel a) shows the conversion efficiency for seed Alfvén waves packets of $cT/\lambda=0.5$. The red line denotes a single AW injected in the northern hemisphere, black and gray lines show conversion efficiencies for seed AWs in both the northern and the southern hemispheres. Panel b) displays conversion efficiencies of AW interactions normalized to the efficiency of a single wave. We show different wavelengths of the injected Alfvén wave (black and blue color) and in-phase versus out-of-phase wave injection in the two hemispheres (thick and thin lines, respectively).
@@
In this section, we explore how the interaction of counter-propagating AWs affects the FMS wave generation during their first emission peak. Single Alfvén waves propagating along curved magnetic field lines can seed fast magnetosonic waves (see Section 2.2, and Yuan et al., 2021; Chen et al., 2024). The efficiency of FMS wave production can change due to non-linear AW interactions (cf. Section 2.1). The first magnetospheric crossing of the seed AWs generates a prompt peak of FMS waves that carries a significant fraction of the total FMS energy. We use the smallest numerical resolution ($N_{R_{\ast}}=128$) that resolves this phase and conduct a broad parameter scan. We ensure sufficiently resolved seed AWs by choosing large injection regions with $\Delta\theta=20^{\circ}$. As a reference for direct comparison, Figure 5 shows the propagation of a single seed AW through the magnetosphere (see animation in the supplementary material, Mahlmann et al., 2024b). Moving along the magnetic field direction, the wave experiences significant shear due to the different lengths of the supporting magnetic field lines (Figure 5, top panels). With increasing shear and relative amplitude of the seed AWs, FMS waves emerge from the inner magnetosphere (bottom panels). The relative amplitude of FMS waves grows as they propagate across the decaying dipole magnetic field (see Equation 9). Electric zones with $E\approx B$ develop due to the propagation effects discussed in Section 2.1 (Figure 5, green contours).
@@
To measure the effect of non-linear wave interactions we inject counter-propagating AWs in both hemispheres. In this scenario (Figure 6), Alfvén waves propagate through the magnetosphere symmetrically and interact in the equatorial region (top panels, see also animation in the supplementary material, out-of-phase: Mahlmann et al., 2024c, in-phase: Mahlmann et al. 2024d). FMS waves are produced and become non-linear in the extended magnetosphere (bottom panels, green contours). To quantify the conversion efficiency, we measure the total Alfvén wave energy injected into the interaction region, $E_{\rm inj}$, by evaluating the energy flux through a sphere close to the stellar surface. We then compare it to the total energy leaving the interaction region as FMS waves, $E_{\rm FMS}$ (measured at some distance from the star). The conversion efficiency is $\eta\equiv E_{\rm FMS}/E_{\rm inj}$. Figure 7 panel a) displays conversion efficiencies for different combinations of seed AWs with wavelength $\lambda=0.5R_{\ast}$ injected for half a wavelength, $cT/\lambda=0.5$. The reference case of a single AW injected in the northern hemisphere with the same properties is shown in red color and can be compared directly to the scaling found by Yuan et al. (2021, dashed line). We probe phase differences of $\varphi=0$ and $\varphi=\pi$ for the interacting half-wavelength modes: the conversion efficiency for in-phase AWs is significantly enhanced.
@@
In Figure 7 panel b), we compute $\eta$ for interacting waves and compare them to those of single waves $\eta_{\rm single}$ varying both envelope size $cT/\lambda\in\left\{0.5,1.0\right\}$ and phase shift $\varphi\in\left\{0,\pi\right\}$. When $\eta/\eta_{\rm single}\lesssim 1$, FMS wave generation is reduced compared to the single wave scenario, while for $\eta/\eta_{\rm single}\gtrsim 1$ the FMS wave production is enhanced. Significant enhancement of the FMS wave generation efficiency happens for short envelopes with $cT/\lambda=0.5$ that are in phase ($\varphi=0$). The enhancement is reduced but not fully suppressed for out-of-phase waves ($\varphi=\pi$), broadly consistent with the analysis of the three-wave interaction coefficient (Section 2.3). Even though the seed AWs are counter-propagating, the conversion coefficient does not vanish due to the shear experienced by waves propagating along dipolar field lines. For increasing wave amplitudes (i.e., larger $b_{\rm eq}$), the conversion efficiency for all configurations decreases compared to that of a single wave. Then, the single wave propagation in a non-uniform (curved) background magnetic field becomes highly nonlinear and dominates over AW interaction for FMS wave generation. The creation of electric zones with rapid dissipation at the collision center (Li et al., 2021) explains the significant suppression of FMS wave production for high-amplitude seed waves with magnetic perturbations that cancel out upon superposition (possible regardless of phase-shift for waves of at least one full wavelength $cT/\lambda=1.0$).
3.4 Repeated interaction of counter-propagating AWs
@@
In this section, we study the enhanced FMS wave generation for short-envelope seed AWs in more detail. With high numerical resolution ($N_{R_{\ast}}=512$), we revisit AWs injected with $cT/\lambda=0.5$, $b_{\rm eq}=0.39$ (see Section 3.3), for different angular extensions $\Delta\theta=5^{\circ}$ and $\Delta\theta=20^{\circ}$. Again, we compare the FMS wave generation during magnetospheric crossings of a single AW with the dynamics of two AWs injected in both hemispheres with phases $\varphi\in\left\{0,\pi\right\}$. Figure 8 shows the time-resolved luminosity carried by escaping FMS waves for injection regions with $\Delta\theta=5^{\circ}$ (bottom panel) and $\Delta\theta=20^{\circ}$ (top panel). AWs without phase difference are the most efficient generators of outgoing FMS waves. Independent of $\Delta\theta$, the energy transported by FMS waves increases by an order of magnitude during a short initial spike. Out-of-phase waves have a reduced conversion efficiency that depends on the width $\Delta\theta$ of the injection site. Waves with large $\Delta\theta$ show a significant decrease in conversion efficiency: approximately two orders of magnitude compared to the phase-aligned case. For the chosen seed waves, we find efficiencies $\eta_{\varphi=0}/\eta_{\rm single}\approx 3.1$, and $\eta_{\varphi=\pi}/\eta_{\rm single}\approx 1.6$, hence, a decrease of approximately $50\%$. Waves with smaller $\Delta\theta$ have a less severe decrease of FMS generation. For the chosen seed waves, we find efficiencies $\eta_{\varphi=0}/\eta_{\rm single}\approx 4.5$, and $\eta_{\varphi=\pi}/\eta_{\rm single}\approx 3.2$, hence, a decrease of approximately $30\%$. Such enhancements were suggested in Section 3.3 and here we confirm them for selected cases with higher numerical resolution.
@@
Figure 8: Outgoing FMS wave luminosity injected by short-envelope AWs. We compare counter-propagating AWs ($b_{\rm eq}\approx 0.39$, $cT/\lambda=0.5$, $\lambda=R_{\ast}/2$, $\theta_{s}=45^{\circ}$) of different phases (black/blue lines) and varying sizes of the injection region (top: $\Delta\theta=20^{\circ}$; bottom: $\Delta\theta=5^{\circ}$) with the FMS wave injected by a single AW (red lines). The inset shows a zoom into the initial interaction period, where the phase difference induces different FMS wave luminosities.
@@
Figure 9: Magnetospheric energy balance with AW absorption at the stellar surface. We vary the absorption coefficient $\mathcal{C}_{\rm abs}$ (see Equation 22). The bottom panel shows the outgoing FMS wave luminosity. The total energy of generated FMS waves decreases at later times. Top panel as in Figure 4, numbers show the amount of absorbed energy.
@@
Besides the strong shear AWs experience when propagating through the magnetosphere, they can be damped plastically when interacting with the magnetar crust (e.g., Li & Beloborodov, 2015). In the last set of tests, we evaluate the dynamics of two AWs injected in opposite hemispheres that experience an energy loss every time they interact with the stellar surface boundary. We adapt the perfectly conducting boundary used throughout this work (see details in Mahlmann & Aloy, 2021; Mahlmann et al., 2023) by including an absorbing layer close to the star. Following Parfrey et al. (2012, equation 53), we add damping terms that drive the field components of the AW polarization ($B^{\phi}$, $E^{r}$, $E^{\theta}$) to the dipole field. We vary the damping rate $\mathcal{C}_{\rm abs}$ for a smooth damping profile active within $r<1.1R_{\ast}$. In practice, we drive the absorption by adding source terms of the form
@@
to the finite volume integration of the respective field components. For the same seed wave properties as above ($b_{\rm eq}\approx 0.39$, $cT/\lambda=0.5$, $\lambda=R_{\ast}/2$, $\theta_{s}=45^{\circ}$, $\Delta\theta=5^{\circ}$), we increase the absorption strength $\mathcal{C}_{\rm abs}$. Figure 9 shows how the magnetospheric excess energy evolves in time (top panel), and how the damping of AWs impacts the outgoing FMS waves (bottom panel). For the explored values of $\mathcal{C}_{\rm abs}$, we compute the relative change of generated FMS wave energy as follows
@@
We then approximate the wave energy absorbed by the boundary layer as the difference $E_{\rm abs}\equiv\Delta E_{\rm tot}-\Delta E_{\rm tot}\left(\mathcal{C}_{\rm abs}=0\right)$, neglecting the small contribution of FMS waves to the total excess energy. The plastic damping of AWs at the stellar boundary only has a small effect on the outgoing FMS waves. While the absorption strength $\mathcal{C}_{\rm abs}$ significantly reduces the total excess energy in the magnetosphere (see Figure 9, top panel), the outgoing FMS wave luminosity decreases only slightly at late times. Specifically, the FMS wave efficiency drops by $0.23\%$ ($\mathcal{C}_{\rm abs}=0.01$), $1.13\%$ ($\mathcal{C}_{\rm abs}=0.05$), $2.06\%$ ($\mathcal{C}_{\rm abs}=0.10$), and $2.83\%$ ($\mathcal{C}_{\rm abs}=0.15$). The initial FMS pulse – emitted before the first boundary interaction of the seed AW – does not experience additional losses. For $t/T_{\rm bounce}[45^{\circ}]\gtrsim 2$, the excess energy (mainly stored in AWs) slowly decays with time. The rate of this late-stage decay is similar for all the probed absorption coefficients $\mathcal{C}_{\rm abs}$. This coincidence of dissipation rates is likely a combination of two effects. First, for the employed resolution ($N_{R_{\ast}}=512$) sheared Alfvén waves could be affected by grid diffusion for $t/T_{\rm bounce}[45^{\circ}]\gtrsim 2$. The chosen resolution $N_{R_{\ast}}$ is a trade-off between computational feasibility and an accurate representation of the relevant physics. Second, as the AWs become increasingly sheared, they could lose their wave character ($\omega=ck_{\parallel}\approx 0$, see also Equation 15) and remain in the inner magnetosphere as an extended shear. Such a state would be less affected by the boundary absorption acting only very close to the star. Figure 9 (top panel) displays the fraction of absorbed energy at different times. We discuss the implications of these wave dynamics in the following sections.

### v2

@@ 3.3 Prompt generation of FMS waves
Figure 5: As Figure LABEL:fig:a1bounce but for a single Alfvén wave injected with $b_{\rm eq}=0.39$ and half a wavelength ($cT/\lambda=0.5$, $\lambda=R_{\ast}/2$) at $\theta_{s}=45^{\circ}$ with width $\Delta\theta=20^{\circ}$ and resolution $N_{r}(N_{R_{\ast}})\times N_{\theta}=19968(512)\times 1600$. The fast wave amplitude relative to the magnetic field can develop extended electrically dominated zones with $E>B$ (green contours, panel d). We provide an animated version of this figure as supplementary material (Mahlmann et al., 2024b). It shows the evolution of the magnetic fields (left animation panel) and electric fields (right animation panel) during times $ct/R_{\ast}=0$ to $16.79$; the real-time duration of the animation is $8\,{\rm s}$.
@@
Figure 6: As Figure 5, but for two Alfvén waves launched symmetrically and out-of-phase ($\varphi=\pi$) in both hemispheres. The Alfvén waves interact in the equatorial region (panels B/C/D) and alter the fast wave generation. We provide an animated version of this figure as supplementary material (Mahlmann et al., 2024c). It shows the evolution of the magnetic fields (left animation panel) and electric fields (right animation panel) during times $ct/R_{\ast}=0$ to $16.79$; the real-time duration of the animation is $8\,{\rm s}$.
@@
Figure 7: Conversion efficiency of injected Alfvén wave energy to outgoing FMS wave energy using $N_{r}(N_{R_{\ast}})\times N_{\theta}=4992(128)\times 400$. Panel a) shows the conversion efficiency for seed Alfvén waves packets of $cT/\lambda=0.5$. The red line denotes a single Alfvén wave injected in the northern hemisphere, black and gray lines show conversion efficiencies for seed Alfvén waves in both the northern and the southern hemispheres. Panel b) displays conversion efficiencies of Alfvén wave interactions normalized to the efficiency of a single wave. We show different wavelengths of the injected Alfvén wave (black and blue color) and in-phase versus out-of-phase wave injection in the two hemispheres (thick and thin lines, respectively).
@@
In this section, we explore how the interaction of counter-propagating Alfvén waves affects the FMS wave generation during their first emission peak. Single Alfvén waves propagating along curved magnetic field lines can seed FMS waves (see Section 2.2, and Yuan et al., 2021; Chen et al., 2024). The efficiency of FMS wave production can change due to nonlinear Alfvén wave interactions (see Section 2.1). The first magnetospheric crossing of the seed Alfvén waves generates a prompt peak of FMS waves that carries a significant fraction of the total FMS energy. We use the smallest numerical resolution ($N_{R_{\ast}}=128$) that resolves this phase and conduct a broad parameter scan. We ensure sufficiently resolved seed Alfvén waves by choosing large injection regions with $\Delta\theta=20^{\circ}$. As a reference for direct comparison, Figure 5 shows the propagation of a single seed Alfvén wave through the magnetosphere (see animation in the supplementary material, Mahlmann et al., 2024b). Moving along the magnetic field direction, the wave experiences significant shear due to the different lengths of the supporting magnetic field lines (Figure 5, top panels). With increasing shear and relative amplitude of the seed Alfvén waves, FMS waves emerge from the inner magnetosphere (bottom panels). The relative amplitude of FMS waves grows as they propagate across the decaying dipole magnetic field (see Equation 9). Electric zones with $E\approx B$ develop due to the propagation effects discussed in Section 2.1 (Figure 5, green contours).
@@
To measure the effect of nonlinear wave interactions we inject counter-propagating Alfvén waves in both hemispheres. In this scenario (Figure 6), Alfvén waves propagate through the magnetosphere symmetrically and interact in the equatorial region (top panels, see also animation in the supplementary material, out-of-phase: Mahlmann et al., 2024c, in-phase: Mahlmann et al. 2024d). FMS waves are produced and become nonlinear in the extended magnetosphere (bottom panels, green contours). To quantify the conversion efficiency, we measure the total Alfvén wave energy injected into the interaction region, $E_{\rm inj}$, by evaluating the energy flux through a sphere close to the stellar surface. We then compare it to the total energy leaving the interaction region as FMS waves, $E_{\rm FMS}$ (measured at some distance from the star). The conversion efficiency is $\eta\equiv E_{\rm FMS}/E_{\rm inj}$. Figure 7 panel a) displays conversion efficiencies for different combinations of seed Alfvén waves with wavelength $\lambda=0.5R_{\ast}$ injected for half a wavelength, $cT/\lambda=0.5$. The reference case of a single Alfvén wave injected in the northern hemisphere with the same properties is shown in red color and can be compared directly to the scaling found by Yuan et al. (2021, dashed line). We probe phase differences of $\varphi=0$ and $\varphi=\pi$ for the interacting half-wavelength modes: the conversion efficiency for in-phase Alfvén waves is significantly enhanced.
@@
In Figure 7 panel b), we compute $\eta$ for interacting waves and compare them to those of single waves $\eta_{\rm single}$ varying both envelope size $cT/\lambda\in\left\{0.5,1.0\right\}$ and phase shift $\varphi\in\left\{0,\pi\right\}$. When $\eta/\eta_{\rm single}\lesssim 1$, FMS wave generation is reduced compared to the single wave scenario, while for $\eta/\eta_{\rm single}\gtrsim 1$ the FMS wave production is enhanced. Significant enhancement of the FMS wave generation efficiency happens for short envelopes with $cT/\lambda=0.5$ that are in-phase ($\varphi=0$). The enhancement is reduced but not fully suppressed for out-of-phase waves ($\varphi=\pi$), broadly consistent with the analysis of the three-wave interaction coefficient (Section 2.3). Even though the seed Alfvén waves are counter-propagating, the conversion coefficient does not vanish due to the shear experienced by waves propagating along dipolar field lines. For increasing wave amplitudes (i.e., larger $b_{\rm eq}$), the conversion efficiency for all configurations decreases compared to that of a single wave. Then, the single wave propagation in a nonuniform (curved) background magnetic field becomes highly nonlinear and dominates over Alfvén wave interaction for FMS wave generation. The creation of electric zones with rapid dissipation at the collision center (Li et al., 2021) explains the significant suppression of FMS wave production for high-amplitude seed waves with magnetic perturbations that cancel out upon superposition (possible regardless of phase shift for waves of at least one full wavelength $cT/\lambda=1.0$).
3.4 Repeated interaction of counter-propagating Alfvén waves
@@
In this section, we study the enhanced FMS wave generation for short-envelope seed Alfvén waves in more detail. With high numerical resolution ($N_{R_{\ast}}=512$), we revisit Alfvén waves injected with $cT/\lambda=0.5$, $b_{\rm eq}=0.39$ (see Section 3.3), for different angular extensions $\Delta\theta=5^{\circ}$ and $\Delta\theta=20^{\circ}$. Again, we compare the FMS wave generation during magnetospheric crossings of a single Alfvén wave with the dynamics of two Alfvén waves injected in both hemispheres with phases $\varphi\in\left\{0,\pi\right\}$. Figure 8 shows the time-resolved luminosity carried by escaping FMS waves for injection regions with $\Delta\theta=5^{\circ}$ (bottom panel) and $\Delta\theta=20^{\circ}$ (top panel). Alfvén waves without phase difference are the most efficient generators of outgoing FMS waves. Independent of $\Delta\theta$, the energy transported by FMS waves increases by an order of magnitude during a short initial spike. Out-of-phase waves have a reduced conversion efficiency that depends on the width $\Delta\theta$ of the injection site. Waves with large $\Delta\theta$ show a significant decrease in conversion efficiency: approximately two orders of magnitude compared to the phase-aligned case. For the chosen seed waves, we find efficiencies $\eta_{\varphi=0}/\eta_{\rm single}\approx 3.1$, and $\eta_{\varphi=\pi}/\eta_{\rm single}\approx 1.6$, hence, a decrease of approximately $50\%$. Waves with smaller $\Delta\theta$ have a less severe decrease of FMS generation. For the chosen seed waves, we find efficiencies $\eta_{\varphi=0}/\eta_{\rm single}\approx 4.5$, and $\eta_{\varphi=\pi}/\eta_{\rm single}\approx 3.2$, hence, a decrease of approximately $30\%$. Such enhancements were suggested in Section 3.3 and here we confirm them for selected cases with higher numerical resolution.
@@
Figure 8: Outgoing FMS wave luminosity injected by short-envelope Alfvén waves. We compare counter-propagating Alfvén waves ($b_{\rm eq}\approx 0.39$, $cT/\lambda=0.5$, $\lambda=R_{\ast}/2$, $\theta_{s}=45^{\circ}$) of different phases (black/blue lines) and varying sizes of the injection region (top: $\Delta\theta=20^{\circ}$; bottom: $\Delta\theta=5^{\circ}$) with the FMS wave injected by a single Alfvén wave (red lines). The inset shows a zoom into the initial interaction period, where the phase difference induces different FMS wave luminosities.
@@
Figure 9: Magnetospheric energy balance with Alfvén wave absorption at the stellar surface. We vary the absorption coefficient $\mathcal{C}_{\rm abs}$ (see Equation 22). The bottom panel shows the outgoing FMS wave luminosity. The total energy of generated FMS waves decreases at later times. Top panel: as in Figure 4, numbers show the amount of absorbed energy.
@@
Besides the strong shear Alfvén waves experience when propagating through the magnetosphere, they can be damped plastically when interacting with the magnetar crust (e.g., Li & Beloborodov, 2015). In the last set of tests, we evaluate the dynamics of two Alfvén waves injected in opposite hemispheres that experience an energy loss every time they interact with the stellar surface boundary. We adapt the perfectly conducting boundary used throughout this work (see details in Mahlmann & Aloy, 2021; Mahlmann et al., 2023) by including an absorbing layer close to the star. Following Parfrey et al. (2012, equation 53), we add damping terms that drive the field components of the Alfvén wave polarization ($B^{\phi}$, $E^{r}$, $E^{\theta}$) to the dipole field. We vary the damping rate $\mathcal{C}_{\rm abs}$ for a smooth damping profile active within $r<1.1R_{\ast}$. In practice, we drive the absorption by adding source terms of the form
@@
to the finite volume integration of the respective field components. For the same seed wave properties as above ($b_{\rm eq}\approx 0.39$, $cT/\lambda=0.5$, $\lambda=R_{\ast}/2$, $\theta_{s}=45^{\circ}$, $\Delta\theta=5^{\circ}$), we increase the absorption strength $\mathcal{C}_{\rm abs}$. Figure 9 shows how the magnetospheric excess energy evolves in time (top panel), and how the damping of Alfvén waves impacts the outgoing FMS waves (bottom panel). For the explored values of $\mathcal{C}_{\rm abs}$, we compute the relative change of generated FMS wave energy as follows:
@@
We then approximate the wave energy absorbed by the boundary layer as the difference $E_{\rm abs}\equiv\Delta E_{\rm tot}-\Delta E_{\rm tot}\left(\mathcal{C}_{\rm abs}=0\right)$, neglecting the small contribution of FMS waves to the total excess energy. The plastic damping of Alfvén waves at the stellar boundary only has a small effect on the outgoing FMS waves. While the absorption strength $\mathcal{C}_{\rm abs}$ significantly reduces the total excess energy in the magnetosphere (see Figure 9, top panel), the outgoing FMS wave luminosity decreases only slightly at late times. Specifically, the FMS wave efficiency drops by $0.23\%$ ($\mathcal{C}_{\rm abs}=0.01$), $1.13\%$ ($\mathcal{C}_{\rm abs}=0.05$), $2.06\%$ ($\mathcal{C}_{\rm abs}=0.10$), and $2.83\%$ ($\mathcal{C}_{\rm abs}=0.15$). The initial FMS pulse – emitted before the first boundary interaction of the seed Alfvén wave – does not experience additional losses. For $t/T_{\rm bounce}[45^{\circ}]\gtrsim 2$, the excess energy (mainly stored in Alfvén waves) slowly decays with time. The rate of this late-stage decay is similar for all the probed absorption coefficients $\mathcal{C}_{\rm abs}$. This coincidence of dissipation rates is likely a combination of two effects. First, for the employed resolution ($N_{R_{\ast}}=512$) sheared Alfvén waves could be affected by grid diffusion for $t/T_{\rm bounce}[45^{\circ}]\gtrsim 2$. The chosen resolution $N_{R_{\ast}}$ is a trade-off between computational feasibility and an accurate representation of the relevant physics. Second, as the Alfvén waves become increasingly sheared, they could lose their wave character ($\omega=ck_{\parallel}\approx 0$, see also Equation 15) and remain in the inner magnetosphere as an extended shear. Such a state would be less affected by the boundary absorption acting only very close to the star. Figure 9 (top panel) displays the fraction of absorbed energy at different times. We discuss the implications of these wave dynamics in the following sections.